OpenStudy (anonymous):

A chemist has one solution that is 80% acid and another solution that is 30% acid. How much of the first (80%) solution is needed to make a 400 L solution that is 62% acid?@Nurali

OpenStudy (anonymous):

.acid + acid = acid 0.80x + 0.30(400-x) = 0.62*400 --------------------- Multiply thru by 100 to get: 80x + 30*400 - 30x = 62*400 50x = 32*400 x = 8*32 x = 256 L (amt. of 80% solution needed in the mixture)